OpenStudy (anonymous):

what is the solution to the system of equations y=2x+5 y+3x=0

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=2x+5 }\) \(\large\color{black}{ \displaystyle y+3x=0 }\) You can do a "substitution". What does the first equation tell you? It tells you that 'y' is the same thing as '2x+5', so you can substitute '2x+5' instead of y, into the other (the second) equation,

OpenStudy (solomonzelman):

and you get: \(\large\color{black}{ \displaystyle \color{blue}{2x+5}+3x=0 }\)

OpenStudy (solomonzelman):

from here, can you solve for \(\large\color{black}{ \displaystyle x}\) in this equation ?